# Janus-faces of Temporal Constraint Languages: A dichotomy of expressivity

### v1

@@ 1. Introduction
1.1. Constraint Satisfaction Problems
The Constraint Satisfaction Problem induced by a relational structure $\mathbb{A}$, denoted by $\operatorname{\textnormal{CSP}}(\mathbb{A})$, is the computational problem of deciding, given a finite input structure $\mathbb{B}$ of the same signature as $\mathbb{A}$, whether there exists a homomorphism from $\mathbb{B}$ to $\mathbb{A}$, i.e. a map preserving all relations. The underlying structure $\mathbb{A}$ will often be referred to as template structure. This notion of fixed-template CSPs provides a uniform framework for modelling many classical computational problems such as graph-colouring problems, $3$-SAT, or solving equations. In 2017, two independent confirmations of a complexity dichotomy for CSPs induced by arbitrary finite structures – conjectured already in [FV93, FV98] – marked a breakthrough in the research programme on the complexity of finite-domain CSPs: for every finite structure $\mathbb{A}$, either $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is solvable in polynomial time, or it is NP-complete [Bul17, Zhu17, Zhu20a], contrasting Ladner’s theorem [Lad75]. Yet, prominent computational problems that can be phrased as the CSPs of a fixed template $\mathbb{A}$ require the domain of $\mathbb{A}$ to be of countably infinite size – for example, the problem of deciding whether a given finite digraph contains a cycle can be phrased as CSP over $(\mathbb{Q};<)$, but cannot be modelled in this way by any finite template. There provably does not exist a complexity dichotomy for CSPs with infinite templates, not even if the template is ‘close to finite’ in the sense of $\omega$-categoricity [BG08, GJK+20]. For certain countably infinite $\omega$-categorical structures, namely first-order reducts of finitely bounded homogeneous structures, though, a P/NP-complete complexity dichotomy has been conjectured by Bodirsky and Pinsker more than a decade ago (see [BPP21, BP16, BKO+17, BKO+19] for various formulations of the conjecture). The conjecture remains wide open in its generality, but has been verified for several classes of structures fitting into the conjectured framework, including temporal constraint languages [BK08b, BK10], equality constraint languages [BK08a], phylogeny CSPs [BJP16, BJP17], the universal homogeneous poset [KP17, KP18], MMSNP [BMM18, BMM21], first-order reducts of any homogeneous undirected graph [BMPP19] including the random graph [BP15a], the universal homogeneous tournament [MP22, MP24], and graph orientation problems with forbidden tournaments [BGP25, BM24, FP25].
@@ 1.2. The algebraic approach
The complexity of finite-domain CSPs is determined by the ‘expressive power’ of the underlying structure $\mathbb{A}$, which is encoded in the set $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of all compatible finitary operations – the polymorphism clone of $\mathbb{A}$. The study of the computational complexity of CSPs via the polymorphisms they have is what is now understood as the algebraic approach to CSPs, first developed in [Jea98, BJK05] and unified by [BOP18]. The finite-domain CSP dichotomy now takes the following form: as the only source of NP-completeness for $\operatorname{\textnormal{CSP}}(\mathbb{A})$ stands the ability of $\mathbb{A}$ to pp-construct EVERYTHING, i.e. every finite structure (see Section 2.1 for the definition of a pp-construction). For the purpose of this paper, we call a structure that pp-constructs EVERYTHING omni-expressive. Every finite structure that lacks this property gives rise to a CSP solvable in polynomial time, and its polymorphism clone satisfies certain symmetries. Here, we say that the polymorphism clone $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of a structure $\mathbb{A}$ satisfies an identity, if the identity is witnessed by operations contained in $\operatorname{\textnormal{Pol}}(\mathbb{A})$ for all evaluations of their arguments. For example, $\operatorname{\textnormal{Pol}}(\mathbb{A})$ always contains
@@
An algebraic approach to CSPs via polymorphisms is for countably infinite structures only possible in the setting of $\omega$-categoricity. Indeed, if $\mathbb{A}$ is $\omega$-categorical, then the complexity of $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is again captured within the ‘local’ algebraic structure of $\operatorname{\textnormal{Pol}}(\mathbb{A})$ [BP15b]. Moreover, for first-order reducts of finitely bounded homogeneous structures $\mathbb{A}$, it is conjectured that omni-expressivity remains the only source of NP-completeness of $\operatorname{\textnormal{CSP}}(\mathbb{A})$ [BPP21, BOP18, BKO+17, BKO+19]. Naturally, the question arises whether the identities (1)-(5) have counterparts for $\omega$-categorical structures that are not omni-expressive. This was answered affirmatively in [BP20] for a pseudo-version of identity (1): every $\omega$-categorical structure that does not pp-construct EVERYTHING has polymorphisms $s,u,$ and $v$ witnessing the identity
@@ 1.3. Loop lemmata
In the early stages of the systematic research programme on CSPs, those induced by finite graphs were among the first to be studied. Observe that the classical $k$-colouring problem coincides with $\operatorname{\textnormal{CSP}}(\mathbb{K}_{k})$, where $\mathbb{K}_{k}$ denotes the clique on $k$ vertices. Vastly generalising the NP-completeness of the $k$-colouring problem for all $k\geq 3$, Hell and Nešetřil showed that every undirected non-bipartite graph is either omni-expressive (and hence its CSP is NP-complete), or has a loop (in which case it is entirely inexpressive, and in particular, its CSP trivial) [HN90]. This result was later extended to certain digraphs:
@@
Let $\mathbb{A}$ be a finite relational structure that is not omni-expressive. If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{G}$-loop condition.
An oligomorphic subgroup $\Omega$ of the automorphism group of a countable digraph $\mathbb{G}$ has, in particular, finitely many orbits in its action on the domain of $\mathbb{G}$. One is thus led to consider the finite quotient of $\mathbb{G}$ modulo $\Omega$, whose vertices are the $\Omega$-orbits, and whose edges are induced from $\mathbb{G}$. Observe that a loop in this finite digraph comes from an edge in $\mathbb{G}$ between two vertices belonging to the same $\Omega$-orbit – a so-called pseudo-loop modulo $\Omega$. The non-existence of pseudo-loops in $\mathbb{G}$ has been identified as a source of computational hardness for $\operatorname{\textnormal{CSP}}(\mathbb{G})$ in several cases [BP20, BBK+23, BKNP25]. Following this direction, we prove a variant of Theorem 1.1 for the case of digraphs $\mathbb{G}$ that are pp-interpretable in a temporal constraint language $\mathbb{A}$ that is not omni-expressive. Roughly, this means that $\mathbb{G}$ is expressible in primitive positive logic over $\mathbb{A}$. Since the structure $(\mathbb{Q};<)$ is $\omega$-categorical, its automorphism group is oligomorphic. The latter naturally acts on $\mathbb{A}$ by automorphisms, and in fact also on $\mathbb{G}$ by the definition of a (pp-)interpretation. Hence, we may regard $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$ also as an oligomorphic subgroup of $\operatorname{\textnormal{Aut}}(\mathbb{G})$.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{G}$ is any smooth digraph that is pp-interpretable in $\mathbb{A}$ and has pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$, then $\mathbb{G}$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{G}$-pseudo-loop condition.
@@
As it turns out, sets of identities of the form as (3) and (4) are better suited for deriving further structural properties that may, in particular, be used in algorithms [BK09, BK14, Bul06, Zhu17, Zhu20a, MP22, MP24, KTV24, PRSS25]. They are formalised through the notion of a $\mathbb{T}$-loop condition for hypergraphs $\mathbb{T}$ of arity $n\geq 3$, defined as a set of n1n-1 identities, explained in detail e.g. in [GJP19]. The following loop lemma for finite symmetric hypergraphs of suitable arity implies – in a similar way to how 1.2 is derived from Theorem 1.1 – the existence of WNU polymorphisms in all not omni-expressive finite structures. Here, a loop in an $n$-ary relation $R$ is a constant tuple $(a.\dots,a)\in R$.
Theorem 1.5 (follows from [BK12], see also [Zhu20b]).
Let $\mathbb{T}$ be a finite symmetric hypergraph of arity $n$ such that $p\nmid n$ for all prime numbers $p\leq|T|$. Either $\mathbb{T}$ is omni-expressive, or $\mathbb{T}$ contains a loop.
Extending the approach used in the proof of Theorem 1.3 to higher arities, we proof a version of Theorem 1.5 for the case symmetric hypergraphs pp-interpretable in a not omni-expressive temporal template. Note that in contrast to the finite version, no number-theoretic restrictions on the arity of $\mathbb{T}$ apply.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{T}$ is any symmetric hypergraph of arity $n\geq 3$ that is pp-interpretable in $\mathbb{A}$, then $\mathbb{T}$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$.
The pseudo-loop condition induced by a hypergraph is again obtained by composing all identities of the corresponding loop-condition with unary function symbols. Since the hypergraphs inducing the Olšák- and the WNU-identities, respectively, are symmetric, the existence of pseudo-Olšák polymorphisms and pseudo-WNU polymorphisms of all arities $n\geq 3$ follows from Theorem 1.6 in the standard way.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{T}$ is any finite symmetric hypergraph of arity $n\geq 3$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{T}$-pseudo-loop condition.
@@ 2.1. Model-theoretic
For $n\in\mathbb{N}$, we denote the set $\{1,\dots,n\}$ by $[n]$. For an $n$-tuple $a=(a_{1},\dots,a_{n})$, we write $a_{i}$ or $\textnormal{pr}_{i}(a)$ for its $i$-th coordinate. By $\ker(a)$ we denote the subset of $n]\times[n]$ consisting of all pairs $(i,j)$ for which $a_{i}=a_{j}$. A relation $R$ on a set $A$ is a subset $R\subseteq A^{n}$ for some $n\in\mathbb{N}$, which is referred to as the arity of $R$. For a subset $I$ of $[n]$, we denote by $\textnormal{pr}_{I}(R)$ the $|I|$-ary relation obtained by projecting $R$ to all of its coordinates $i$ with $i\in I$. We also write $R|_{B_{1}\times\dots\times B_{n}}$ for the relation $R\cap(B_{1}\times\dots\times B_{n})$, where $B_{i}\subseteq A$. For the purpose of this paper, we assume all relations considered to be non-empty. A (relational) structure is a tuple $\mathbb{A}=(A;\mathcal{R})$ consisting of a set $A$ and a finite family $\mathcal{R}$ of relations on $A$ indexed by relational symbols. For structures $\mathbb{A}=(A;\mathcal{R})$ and $\mathbb{B}=(B;\mathcal{R}^{\prime})$ indexed the same relational symbols, a homomorphism from $\mathbb{A}$ to $\mathbb{B}$ is a mapping $f:A\to B$ such that for every relation $R\in\mathcal{R}$ and every $(r_{1},\dots,r_{n})\in R$, the tuple $(f(r_{1}),\dots,f(r_{n}))$ is contained in the corresponding relation of $\mathcal{R}^{\prime}$. The structures $\mathbb{A}$ and $\mathbb{B}$ are homomorphically equivalent if there exists homomorphisms both from $\mathbb{A}$ to $\mathbb{B}$, and from $\mathbb{B}$ to $\mathbb{A}$. An automorphism of $\mathbb{A}$ is a bijective homomorphism $f:A\to A$ whose inverse mapping $f^{-1}$ is also a homomorphism. By $\operatorname{\textnormal{Aut}}(\mathbb{A})$ we denote the automorphism group of $\mathbb{A}$. In the following, by componentwise application we will understand all mappings also as functions on $n$-tuples.
@@
A first-order formula $\phi$ is called primitive positive (pp) over a set consisting of relational symbols $R_{1},\dots,R_{n}$ if it involves only the predicates $R_{i}$, existential quantification, and conjunction. A relation is pp-definable in a structure $\mathbb{A}=(A;\mathcal{R})$ if it is definable by a pp-formula over the relational symbols from $\mathcal{R}$. A structure $\mathbb{B}$ is pp-definable in $\mathbb{A}$ if all of its relations are. We say that $\mathbb{A}$ $p$-interprets $\mathbb{B}$ if there exist $k\geq 1$ and a partial surjective map $h:A^{k}\to B$ with the property that for every relation $R\subseteq B^{n}$ that is either a relation of $\mathbb{B}$, the equality relation on $B$, or $B$, its preimage $h^{-1}(R)$ – regarded as a relation of arity $nk$ on $A$ – is pp-definable in $\mathbb{A}$. If $\mathbb{A}$ $p$-interprets $\mathbb{B}$, then $\operatorname{\textnormal{CSP}}(\mathbb{B})$ is log-space reducible to $\operatorname{\textnormal{CSP}}(\mathbb{A})$ [BJK05]. We say that $\mathbb{A}$ $p$-constructs $\mathbb{B}$ if $\mathbb{B}$ is homomorphically equivalent to a structure that is pp-interpretable in $\mathbb{A}$. As homomorphically equivalent structures have the same CSPs, this also implies a log-space reduction from $\operatorname{\textnormal{CSP}}(\mathbb{B})$ to $\operatorname{\textnormal{CSP}}(\mathbb{A})$. We say that $\mathbb{A}$ pp-constructs EVERYTHING if $\mathbb{A}$ pp-constructs every finite structure; we then call $\mathbb{A}$ omni-expressive. Clearly, the CSP induced by an omni-expressive structure is NP-complete. By the finite-domain CSP dichotomy [Bul17, Zhu17, Zhu20a], the CSP of a finite structure that is not omni-expressive is always solvable in polynomial time.
@@
A permutation group $\Omega$ acting on a set $A$ induces an equivalence relation on $A$: two elements $a,b\in A$ are equivalent if there is $\alpha\in\Omega$ such that $\alpha(a)=b$. The corresponding equivalence classes are called the orbits of $\Omega\curvearrowright A$, or simply $\Omega$-orbits. By componentwise evaluation, $\Omega$ acts on $A^{k}$ for every $k\in\mathbb{N}.$ The group is oligomorphic if this action has finitely many orbits for every $k\in\mathbb{N}$. A countable structure $\mathbb{A}$ is $\omega$-categorical if $\operatorname{\textnormal{Aut}}(\mathbb{A})$ is oligomorphic.

### v2

@@ 1. Introduction
The Constraint Satisfaction Problem induced by a relational structure $\mathbb{A}$, denoted by $\operatorname{CSP}(\mathbb{A})$, is the computational problem of deciding, given a finite input structure $\mathbb{B}$ of the same signature as $\mathbb{A}$, whether there exists a homomorphism from $\mathbb{B}$ to $\mathbb{A}$, i.e. a map preserving all relations. The underlying structure $\mathbb{A}$ will often be referred to as template structure. This notion of fixed-template CSPs provides a uniform framework for modelling many classical computational problems such as graph-colouring problems, $3$-SAT, or solving equations. In 2017, two independent confirmations of a complexity dichotomy for CSPs induced by arbitrary finite structures – conjectured already in [FV93, FV98] – marked a breakthrough in the research programme on the complexity of finite-domain CSPs: for every finite structure $\mathbb{A}$, either $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time, or it is NP-complete [Bul17, Zhu17, Zhu20a], contrasting Ladner’s theorem [Lad75]. Yet, prominent computational problems that can be phrased as the CSPs of a fixed template $\mathbb{A}$ require the domain of $\mathbb{A}$ to be of countably infinite size – for example, the problem of deciding whether a given finite digraph contains a cycle can be phrased as CSP over $(\mathbb{Q};<)$, but cannot be modelled in this way by any finite template. There provably does not exist a complexity dichotomy for CSPs with infinite templates, not even if the template is ‘close to finite’ in the sense of $\omega$-categoricity [BG08, GJK+20]. For certain countably infinite $\omega$-categorical structures, namely first-order reducts of finitely bounded homogeneous structures, though, a P/NP-complete complexity dichotomy has been conjectured by Bodirsky and Pinsker more than a decade ago (see [BPP21, BP16, BKO+17, BKO+19] for various formulations of the conjecture). The conjecture remains wide open in its generality, but has been verified for several classes of structures fitting into the conjectured framework, including temporal constraint languages [BK08b, BK10], equality constraint languages [BK08a], phylogeny CSPs [BJP16, BJP17], the universal homogeneous poset [KP17, KP18], MMSNP [BMM18, BMM21], first-order reducts of any homogeneous undirected graph [BMPP19] including the random graph [BP15a], the universal homogeneous tournament [MP22, MP24], and graph orientation problems with forbidden tournaments [BGP25, BM24, FP25].
@@ 1.1. The algebraic approach
The complexity of finite-domain CSPs is determined by the ‘expressive power’ of the underlying structure $\mathbb{A}$, which is encoded in the set $\operatorname{Pol}(\mathbb{A})$ of all compatible finitary operations – the polymorphism clone of $\mathbb{A}$. The study of the computational complexity of CSPs via the polymorphisms they have is what is now understood as the algebraic approach to CSPs, first developed in [Jea98, BJK05] and unified by [BOP18]. The finite-domain CSP dichotomy now takes the following form: as the only source of NP-completeness for $\operatorname{CSP}(\mathbb{A})$ stands the ability of $\mathbb{A}$ to pp-construct EVERYTHING, i.e. every finite structure (see Section 2.1 for the definition of a pp-construction). For the purpose of this paper, we call a structure that pp-constructs EVERYTHING omni-expressive. Every finite structure that lacks this property gives rise to a CSP solvable in polynomial time, and its polymorphism clone satisfies certain symmetries. Here, we say that the polymorphism clone $\operatorname{Pol}(\mathbb{A})$ of a structure $\mathbb{A}$ satisfies an identity, if the identity is witnessed by operations contained in $\operatorname{Pol}(\mathbb{A})$ for all evaluations of their arguments. For example, $\operatorname{Pol}(\mathbb{A})$ always contains
@@
An algebraic approach to CSPs via polymorphisms is for countably infinite structures only possible in the setting of $\omega$-categoricity. Indeed, if $\mathbb{A}$ is $\omega$-categorical, then the complexity of $\operatorname{CSP}(\mathbb{A})$ is again captured within the ‘local’ algebraic structure of $\operatorname{Pol}(\mathbb{A})$ [BP15b]. Moreover, for first-order reducts of finitely bounded homogeneous structures $\mathbb{A}$, it is conjectured that omni-expressivity remains the only source of NP-completeness of $\operatorname{CSP}(\mathbb{A})$ [BPP21, BOP18, BKO+17, BKO+19]. Naturally, the question arises whether the identities (1)-(5) have counterparts for $\omega$-categorical structures that are not omni-expressive. This was answered affirmatively in [BP20] for a pseudo-version of identity (1): every $\omega$-categorical structure that does not pp-construct EVERYTHING has polymorphisms $s,u,$ and $v$ witnessing the identity
@@ 1.2. Loop lemmata
In the early stages of the systematic research programme on CSPs, those induced by finite graphs were among the first to be studied. Observe that the classical $k$-colouring problem coincides with $\operatorname{CSP}(\mathbb{K}_{k})$, where $\mathbb{K}_{k}$ denotes the clique on $k$ vertices. Vastly generalising the NP-completeness of the $k$-colouring problem for all $k\geq 3$, Hell and Nešetřil showed that every undirected non-bipartite graph is either omni-expressive (and hence its CSP is NP-complete), or has a loop (in which case it is entirely inexpressive, and in particular, its CSP trivial) [HN90]. This result was later extended to certain digraphs:
@@
Let $\mathbb{A}$ be a finite relational structure that is not omni-expressive. If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{Pol}(\mathbb{A})$ satisfies the $\mathbb{G}$-loop condition.
1.3. Our contribution
An oligomorphic subgroup $\Omega$ of the automorphism group of a countable digraph $\mathbb{G}$ has, in particular, finitely many orbits in its action on the domain of $\mathbb{G}$. One is thus led to consider the finite quotient of $\mathbb{G}$ modulo $\Omega$, whose vertices are the $\Omega$-orbits, and whose edges are induced from $\mathbb{G}$. Observe that a loop in this finite digraph comes from an edge in $\mathbb{G}$ between two vertices belonging to the same $\Omega$-orbit – a so-called pseudo-loop modulo $\Omega$. The non-existence of pseudo-loops in $\mathbb{G}$ has been identified as a source of computational hardness for $\operatorname{CSP}(\mathbb{G})$ in several cases [BP20, BBK+23, BKNP25]. Following this direction, we prove a variant of Theorem 1.1 for the case of digraphs $\mathbb{G}$ that are pp-interpretable in a temporal constraint language $\mathbb{A}$ that is not omni-expressive. Roughly, this means that $\mathbb{G}$ is expressible in primitive positive logic over $\mathbb{A}$. Since the structure $(\mathbb{Q};<)$ is $\omega$-categorical, its automorphism group is oligomorphic. The latter naturally acts on $\mathbb{A}$ by automorphisms, and in fact also on $\mathbb{G}$ by the definition of a (pp-)interpretation. Hence, we may regard $\operatorname{Aut}((\mathbb{Q};<))$ also as an oligomorphic subgroup of $\operatorname{Aut}(\mathbb{G})$.
@@ 1.3. Our contribution
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{G}$ is any smooth digraph that is pp-interpretable in $\mathbb{A}$ and has pseudo-algebraic length $1$ modulo $\operatorname{Aut}((\mathbb{Q};<))$, then $\mathbb{G}$ contains a pseudo-loop modulo $\operatorname{Aut}((\mathbb{Q};<))$.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{Pol}(\mathbb{A})$ satisfies the $\mathbb{G}$-pseudo-loop condition.
@@
As it turns out, sets of identities of the form as (3) and (4) are better suited for deriving further structural properties that may, in particular, be used in algorithms [BK09, BK14, Bul06, Zhu17, Zhu20a, MP22, MP24, KTV24, PRSS25]. They are formalised through the notion of a $\mathbb{T}$-loop condition for hypergraphs $\mathbb{T}$ of arity $n\geq 3$, defined as a set of n1n-1 identities, as explained in detail e.g. in [GJP19].
Extending our approach to higher arities, we prove a version of Theorem 1.3 for the case of hypergraphs pp-interpretable in a not omni-expressive temporal template. A pseudo-loop in an $n$-ary hypergraph is a hyperedge all of whose coordinates belong to the same orbit; a hypergraph is cyclic if its hyperedges are invariant under cyclic permutations, and it is $2$-transitive if it is invariant under the action of some $2$-transitive subgroup of the symmetric group on $n$ elements.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{T}$ is any cyclic and $2$-transitive hypergraph of arity $n\geq 3$ that is pp-interpretable in $\mathbb{A}$, then $\mathbb{T}$ contains a pseudo-loop modulo $\operatorname{Aut}((\mathbb{Q};<))$.
The pseudo-loop condition induced by a hypergraph is again obtained by composing all identities of the corresponding loop-condition with unary function symbols. Since the hyperedges of the hypergraphs inducing the Olšák- and the WNU-identities are invariant under all permutations, they satisfy in particular the imposed symmetry conditions; hence, the existence of pseudo-Olšák polymorphisms and pseudo-WNU polymorphisms of all arities $n\geq 3$ follows from Theorem 1.5 in the standard way.
@@
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{T}$ is any finite cyclic and $2$-transitive hypergraph of arity $n\geq 3$, then $\operatorname{Pol}(\mathbb{A})$ satisfies the $\mathbb{T}$-pseudo-loop condition.
1.4. Is $2$-transitivity the new cyclicity?
Let us remark that the conclusion of Theorem 1.5 clearly fails in the case $n=2$, as it would amount to the existence of a binary polymorphism that is pseudo-commutative – yet temporal constraint languages preserved by the binary operations $\ell\ell$ or $\operatorname{mi}$ are neither omni-expressive, nor must they admit such polymorphisms. It might be of interest to point out that pseudo-loop lemmata involving symmetry conditions provided by non-trivial group actions have also been established for finite hypergraphs of arity $n\geq 3$ (see [BK12, Zhu20b, Bru23, Zhu24]). Most influentially, the study of finite cyclic hypergraphs in [BK12] stands as a landmark result in this context. In the temporal setting, cyclicity by itself is not a sufficiently strong condition: as mentioned before, there exist not-omni expressive temporal constraint languages that do not admit any pseudo-cyclic polymorphisms. A concrete example is the structure $(\mathbb{Q};<,x\neq y\vee u=v)$ [Bod21, Proposition], highlighting the necessity of the $2$-transitivity condition on top of the cyclicity condition in Theorem 1.5. Observe that these symmetry conditions are proper in the sense that a hypergraph may be cyclic and $2$-transitive without being fully symmetric, i.e. invariant under all permutations of its hyperedges. For example, for odd $n\geq 5$ the alternating group $A_{n}$ is a proper 2-transitive subgroup of the full symmetric group $S_{n}$ which contains all cyclic permutations.
Whether or not cyclicity is truly a necessary condition in Theorem 1.5 remains an interesting open question for future work. For hypergraphs of arity $n=3$, it is readily implied by 2-transitivity, whereas for all arities $n>3$, it is not.
@@ 2.1. Model-theoretic
For $n\in\mathbb{N}$, we denote the set $\{1,\dots,n\}$ by $[n]$. For an $n$-tuple $a=(a_{1},\dots,a_{n})$, we write $a_{i}$ or $\operatorname{pr}_{i}(a)$ for its $i$-th coordinate. By $\ker(a)$ we denote the subset of $n]\times[n]$ consisting of all pairs $(i,j)$ for which $a_{i}=a_{j}$. A relation $R$ on a set $A$ is a subset $R\subseteq A^{n}$ for some $n\in\mathbb{N}$, which is referred to as the arity of $R$. For a subset $I$ of $[n]$, we denote by $\operatorname{pr}_{I}(R)$ the $|I|$-ary relation obtained by projecting $R$ to all of its coordinates $i$ with $i\in I$. For the purpose of this paper, we assume all relations considered to be non-empty. A (relational) structure is a tuple $\mathbb{A}=(A;\mathcal{R})$ consisting of a set $A$ and a finite family $\mathcal{R}$ of relations on $A$ indexed by relational symbols. For structures $\mathbb{A}=(A;\mathcal{R})$ and $\mathbb{B}=(B;\mathcal{R}^{\prime})$ indexed the same relational symbols, a homomorphism from $\mathbb{A}$ to $\mathbb{B}$ is a mapping $f:A\to B$ such that for every relation $R\in\mathcal{R}$ and every $(r_{1},\dots,r_{n})\in R$, the tuple $(f(r_{1}),\dots,f(r_{n}))$ is contained in the corresponding relation of $\mathcal{R}^{\prime}$. The structures $\mathbb{A}$ and $\mathbb{B}$ are homomorphically equivalent if there exists homomorphisms both from $\mathbb{A}$ to $\mathbb{B}$, and from $\mathbb{B}$ to $\mathbb{A}$. An automorphism of $\mathbb{A}$ is a bijective homomorphism $f:A\to A$ whose inverse mapping $f^{-1}$ is also a homomorphism. By $\operatorname{Aut}(\mathbb{A})$ we denote the automorphism group of $\mathbb{A}$. In the following, by componentwise application we will understand all mappings also as functions [Zhu24] on $n$-tuples.
@@
A first-order formula $\phi$ is called primitive positive (pp) over a set consisting of relational symbols $R_{1},\dots,R_{n}$ if it involves only the predicates $R_{i}$, existential quantification, and conjunction. A relation is pp-definable in a structure $\mathbb{A}=(A;\mathcal{R})$ if it is definable by a pp-formula over the relational symbols from $\mathcal{R}$. A structure $\mathbb{B}$ is pp-definable in $\mathbb{A}$ if all of its relations are. We say that $\mathbb{A}$ $p$-interprets $\mathbb{B}$ if there exist $k\geq 1$ and a partial surjective map $h:A^{k}\to B$ with the property that for every relation $R\subseteq B^{n}$ that is either a relation of $\mathbb{B}$, the equality relation on $B$, or $B$, its preimage $h^{-1}(R)$ – regarded as a relation of arity $nk$ on $A$ – is pp-definable in $\mathbb{A}$. If $\mathbb{A}$ $p$-interprets $\mathbb{B}$, then $\operatorname{CSP}(\mathbb{B})$ is log-space reducible to $\operatorname{CSP}(\mathbb{A})$ [BJK05]. We say that $\mathbb{A}$ $p$-constructs $\mathbb{B}$ if $\mathbb{B}$ is homomorphically equivalent to a structure that is pp-interpretable in $\mathbb{A}$. As homomorphically equivalent structures have the same CSPs, this also implies a log-space reduction from $\operatorname{CSP}(\mathbb{B})$ to $\operatorname{CSP}(\mathbb{A})$. We say that $\mathbb{A}$ pp-constructs EVERYTHING if $\mathbb{A}$ pp-constructs every finite structure; we then call $\mathbb{A}$ omni-expressive. Clearly, the CSP induced by an omni-expressive structure is NP-complete. By the finite-domain CSP dichotomy [Bul17, Zhu17, Zhu20a], the CSP of a finite structure that is not omni-expressive is always solvable in polynomial time.
@@
A permutation group $\Omega$ acting on a set $A$ induces an equivalence relation on $A$: two elements $a,b\in A$ are equivalent if there is $\alpha\in\Omega$ such that $\alpha(a)=b$. The corresponding equivalence classes are called the orbits of $\Omega\curvearrowright A$, or simply $\Omega$-orbits. By componentwise evaluation, $\Omega$ acts on $A^{k}$ for every $k\in\mathbb{N}.$ The group is oligomorphic if this action has finitely many orbits for every $k\in\mathbb{N}$. A countable structure $\mathbb{A}$ is $\omega$-categorical if $\operatorname{Aut}(\mathbb{A})$ is oligomorphic.